The volume inside the cylinders of all of the pistons in a combustion engine is typically specified in cc (cm3). A typical 4-cylinder car engine has a cylinder volume of 3900. cm3. Immediately after ignition, the pressure in the cylinder is approximately 125 atm and the temperature is 320. °C. Assuming the gas mixture in the cylinder behaves ideally, and is occupying the maximum volume of the cylinders, how many moles of gas are in the cylinders immediately after ignition?
Answer #1

total volume of four cylinders,V = (3900 x 4) cc = 15600 cc

since

1 cc(cm3) = 1 mL

hence

15600 cc = 15600 mL = 15.6 L

Temperature,T= 320 ° C = (320 + 273.15) K = 593.15 K

Pressure, P = 125 atm

by the ideal gas equeation

PV = nRT

number of moles, n = PV/RT

= 125 * 15.6 / ( 0.0821 * 593.15 )

  n = 40.04 moles